OpenStudy (anonymous):

What is the solution set for |–2x + 8| = 12 Answer x = 2 and x = 10 x = –2 and x = 10 x = –2 and x = –10 x = 2 and x = –10

OpenStudy (amistre64):

what do you get when you plug in the options?

OpenStudy (anonymous):

its the first one right?

OpenStudy (amistre64):

i dont think it would be, but we can do better than just guess

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

what is -2 times 2?

OpenStudy (anonymous):

-4

OpenStudy (amistre64):

and -4 + 8 = ?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

would you say that |4| = 12 then ?

OpenStudy (anonymous):

no...

OpenStudy (amistre64):

then x cannot equal 2

OpenStudy (amistre64):

lets try 10 :)

OpenStudy (amistre64):

what is -2 times 10?

OpenStudy (anonymous):

-20

OpenStudy (amistre64):

and -20 + 8 = ?

OpenStudy (anonymous):

12

OpenStudy (amistre64):

well, -12 yes; does |-12| = 12 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x = –2 and x = 10

OpenStudy (amistre64):

then we conclude that x does not equal 2; and it does equal 10

OpenStudy (anonymous):

amistre can u stay here? im doing attest

OpenStudy (anonymous):

A test and I'm gonna need some more help

OpenStudy (amistre64):

hmmm ..... doing a test doesnt sound very ethical to me; I can help, but I want you to find the answers

OpenStudy (anonymous):

I knwo I just need you to help me figure out how to solve the problems and there is only 10 more questions

OpenStudy (amistre64):

lets do the next one then and see where it leads

OpenStudy (anonymous):

What are the solution intervals for |x + 5| > 3? Answer x < –8 or x > –2 x > –8 or x < –2 x < 8 or x > –2 x > 8 or x < –2

OpenStudy (amistre64):

lets start out by seeing of x=0 is good or bad

OpenStudy (anonymous):

wouldnt it be good?

OpenStudy (amistre64):

when x=0 we get: |5|>3 is this a true or false result?

OpenStudy (anonymous):

true

OpenStudy (amistre64):

it looks good to me; so we can narrow it down to between 2 numbers

OpenStudy (amistre64):

\[\cancel{x < –8 or x > –2}\] x > –8 or x < –2 x < 8 or x > –2 \[\cancel{x > 8 or x < –2}\]

OpenStudy (amistre64):

lets try x = 7

OpenStudy (anonymous):

it works

OpenStudy (amistre64):

then this one looks to be the right area x < 8 or x > –2

OpenStudy (amistre64):



OpenStudy (anonymous):

Ok sweet~ here is the next one On the graph of the equation 3x – 2y = –18, what is the value of the y-intercept? Answer 6 –9 9 –6

OpenStudy (amistre64):

when x=0; what does y=?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

or 1?

OpenStudy (amistre64):

i would say neither since 0 or 1 is not an option is it

OpenStudy (anonymous):

ugh this is confusing

OpenStudy (amistre64):

lets try it again: when x = 0; y = ? 3x – 2y = –18 0 ------------ 0 -2y = -18

OpenStudy (anonymous):

-9?

OpenStudy (anonymous):

or actually its 9

OpenStudy (amistre64):

close, but two negatives give a + result; id go 9 :)

OpenStudy (anonymous):

so the answer is 9 right?

OpenStudy (amistre64):

i would go with 9 yes

OpenStudy (anonymous):

What is the slope of the line which passes through (–6, 0) and (4, 0)? Answer 0 –2 undefined 2

OpenStudy (anonymous):

dont you do like 0-0 and 4-(-6) or something liek that>

OpenStudy (amistre64):

both your y values are the same; which means the no matter what you pick for x; you always have the same y value

OpenStudy (anonymous):

like y2-y1 and x2-x1